

********************************************************************
UARS WEEKLY STATUS REPORT - 27 March 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data this week without apparent problems.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE was powered off at the end of day on March 25, 1998. All engineering
parameters were nominal prior to turning off and the science data products
continue to be of excellent quality.
***********
MLS
***********
MLS was reactivated successfully on Mar 26, in the reduced 85 watt power mode.
MLS instrument operations have continued flawlessly. MLS is scheduled to
operate through Mar 29.
***********
HRDI
***********
HRDI is operating. Scheduled shutdown is March 27. Next turn on will be
April 5.
***********
WINDII
***********
The following papers were presented during March 16-20, 1998 at the
International Symposium on Dynamics and Structure of the Mesopause
Region (DYSMER), held at the Uji campus of Kyoto University, and
organized by the Radio Atmospheric Science Center, with co-sponsorship
by a number of organizations.
1. M.G. Shepherd, C. Gardner, B. Prawirosoehardjo, S.-P. Zhang and B. Solheim
Planetary scale and tidal perturbations in mesopheric temperature
observed by WINDII
2. W.E. Ward
Energy and constituent budgets near the mesopause: the implication
of observations of large scale variability in the airglow
3. G.G. Shepherd, S.-P. Zhang and D.-Y. Wang
Variability in mnesosphere/lower thermosphere dynamics and
species concentrations observed by WINDII
There were a number of other papers using or making reference to
HRDI, WINDII and other UARS data. Thus while the coincidence with
the Science Team meeting was unfortunate, the significant contributions
of UARS to mesospheric science were clearly recognized, and the overall
result was very beneficial to the community.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF performed the UARS facility power down and power up for the
building 23 power outage during the 03/20/98 - 03/23/98 period. No significant
problems occurred during the power up. The cluster was up and operational at
approximately 09:30AM local time.
The CDHF implemented CMS-RAC-MPG communications with Multinet during the
outage period. The CDHF worked with CMS, MPG and RAC Instrumentors to
communicate directly with MPG and CMS in Building 32, without routing through
Building 23. Also, all RAC transfers were successfully transmitted prior to
system power down.
The CDHF has received a total of 380 CD-ROM requests as of 27-March-1998.
There was one order received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
20 days PEM (3.32) PROD (1379R-1049C=330L)
84 days MLS (4.90) PROD (1297R-1139C-17F=141L)
9 days CLAES (7.0) ENG (38R-33C=5L)
0 days HAL (4.80) ENG (4R-4C=0L)
The CDHF loaded January 1998 current day DAAC data.
Pending software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 March 1998 (Orbit 35540, GMT Day 073) through 20
March 1998 (Orbit 35644, GMT Day 079).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 074/0544 35558 Off
HALOE 074/0722 35559 On
HALOE 075/1505 35579 Off
HALOE 075/1649 35580 On
MLS 075/2333 35584 Off
ACRIM 076/0046 35585 On
HRDI 076/0046 35585 On
SUSIM 076/1321 35593 On
The beta angle reached a maximum angle of 54.5 degrees on 15 March 1998
(DOY 074). The beta angle is now increasing toward a minimum angle of 0.0
degrees on 02 April 1998 (DOY 092).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.3 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 14 March 1998 and 19 March 1998 are provided in Appendix D.
The clock error ranged from -11.1 to +10.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% ACRIM 01
% HALOE 06
% HRDI 01
% MLS 01
% SOLSTICE 01
% SUSIM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
Two (2) Anomaly Reports remained OPEN at the end of this report period.
TRANSPONDER FREQUENCIES:
Transponder A: 98/041 - 2287.497104
Transponder B: 98/042 - 2287.499330
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/078 show the observatory data loss to be
20 hours, 42 minutes, 16.660 seconds (an increase of 8.576 seconds). This
is a 0.0364 percent data loss which equals a 99.9636 percent data capture
for the mission.
The increase of 8.576 seconds occurred over two days as follows:
7 A 0.224 second loss on Day 073 due to a hit in the transmission line, and
7 A 8.352 second loss on Day 077 due to a MSOCC Telemetry And Command
(TAC) computer crash.
Building 23 power was removed as scheduled on Friday, 20 March 1998. No
problems with the backup planning have been observed.
The FOT supported CNMOS year 2000 testing on 20 March 1998. The test was
simple and functioned as a quick Rsmoke testS of the systems. The modified
Application Processor (AP) system software with the year set to 2000 was
used in a UARS data flow configuration. UTTS data was input to the AP and
the data appeared satisfactory. A command was transmitted and a data file
was pulled from the CMS. Additional and more detail testing is planned.
No IP testing was executed by the FOT this week. The FOT plans on
participating in the forward link IP testing beginning next week.
The FOT is still unable to receive good Q-channel OBC dump data.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. File transfer failures (between CDHF and UMPGAL) occurred 3 days
last week. The relatively high frequency of file transfer failures between
UMPGAL and CDHF has probably been due to the numerous reconfigurations of
the network communication links. This problem is likely to continue until
the UMPGAL cluster and network interface configurations are finalized (at
least one month). Given the amount of manual effort required to restore
parallel level-0 processing when the file transfer link fails, it is not
practical to continue the parallel operating mode until these configuration
changes are complete. Therefore, all operational processing of level-0
telemetry files on UMPGAL will cease until further notice.
The FOT has received (20 March) a request from the NCC Database manager for
permission to delete several UARS SUPIDENs from the NCC database. The FOT
will investigate whether or not the identified SUPIDENs are still necessary
for UARS operations before permission for the deletions is given.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
Software development is now tentatively completed, and the software is
currently being tested with the UTTS. Testing was performed successfully
on Thursday and Friday (19 and 20 March 1998), and all tested features
worked as expected. A new database will now be generated to support the
software, and final testing will be completed when the new database is in
place. Generation of the new database is contingent upon operational
release of the new AP software, which is anticipated this week.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 02 April 1998.
A second building 23 power outage is scheduled from about 12 pm, 24 April
1998, until about 12 pm, 27 April 1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 20 March 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35527 072/0255 98-094 WINDII Power ON
35537 072/1958 98-095 HALOE Power OFF
35539 072/2319 98-096 HALOE Power ON
35558 074/0544 98-097 HALOE Power OFF
35559 074/0722 98-098 HALOE Power ON
35579 075/1505 98-099 HALOE Power OFF
35580 075/1649 98-100 HALOE Power ON
35584 075/2333 98-101 SOLSTICE Power ON
35584 075/2333 98-102 MLS Power OFF
35585 076/0046 98-103 ACRIM Power ON
35585 076/0046 98-104 HRDI Power ON
35593 076/1321 98-105 SUSIM Power ON
35593 076/1321 98-106 OBC RTS 12 reload
35597 076/2013 98-107 OBC Clk Adj fr 20.6 to 0.00 msec/day
35603 077/0606 98-108 SUSIM Adjustment
35618 078/0610 98-109 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 20 March 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 35526-77,85-99
SORLBXTMPR Y-LO Logic_Box Temperature 35526-77
HATBAT Y-HI Bolometer Area Temp 35535-6,78
ACTEMPANA1 Y-LO Temp PCU Board 35592-600
ACTEMPANA2 Y-LO Temp Sensor Electronics 35592-97
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 20 March 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35597 076/20:21 (17 Mar 98) from 20.6 to 00.0 msec/day
35618 078/06:18 (19 Mar 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 March 1998
14 Mar 1998 - GMT Day 073 Beta = 54.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +56.0/ -11.2 +5.79/+4.51 +25.1/-16.0 25.1 1.020 21.9
3 +84.0/ -11.2 +4.51/+3.23 +24.2/-16.0 25.1 1.022 21.4
19 Mar 1998 - GMT Day 078 Beta = 48.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.19/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +33.6/ -44.8 +5.79/+4.51 +26.5/-15.7 25.4 1.025 21.9
3 +89.6/ -11.2 +4.51/+3.23 +25.8/-15.8 25.3 1.026 21.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 20 March 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-011 FOT 35533 072/1314 N/A TAC Crashed
98-009 FOT 35585 076/0046 14170 Generic Late Acq
98-012 FOT 35595 076/1622 19757 Late Acq
98-013 FOT 35603 077/0606 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-007 FOT Feb 1998 n/a 14170 Generic Late Acq
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period